# Reservoir Computing as a Language Model

### v1

@@ I Introduction
Input Textshak$\begin{pmatrix}0.2\\ \vdots\\ 0.8\end{pmatrix}$$\begin{pmatrix}0.3\\ \vdots\\ 0.9\end{pmatrix}$$\begin{pmatrix}0.5\\ \vdots\\ 0.6\end{pmatrix}$$\begin{pmatrix}0.7\\ \vdots\\ 0.1\end{pmatrix}$TimeReservoirAtt-Enhanced ReservoirTransformer$\begin{pmatrix}w_{1}\\ w_{2}\\ \vdots\\ w_{m}\end{pmatrix}$ Output Layer$\begin{pmatrix}p_{a}=0.1\\ p_{b}=0.01\\ p_{c}=0.12\\ \vdots\end{pmatrix}$ $H(y,\hat{y})=-\sum_{i}y_{i}\log\hat{y}_{i}$ Cross Entropy – PredictionVector Embedding
@@ II Task Formulation and Dataset
We tackle a character-level next-character prediction task. Let $\mathbf{X}=\{x_{1},\dots,x_{T}\}$ represent a fixed-length input sequence of characters, and let $y_{T+1}$ be the next character to predict. We treat the problem as a multi-class classification over the vocabulary $\mathcal{V}=59$. The dataset is a 9-million-character corpus of Shakespeare text inspired by recent minimalist transformer implementations such as Karpathy’s NanoGPT [22], which we split into $6$ shards: $5$ shards for training and $1$ shard for testing. Each shard is processed at the character level, forming sequences of length $32$ for the input, plus the next character as the target. Before training the text corpus is preprocessed by lowercasing all letters to achieve a smaller vocabulary $\mathcal{V}=59$. Training optimizes a standard cross-entropy loss over the predicted logits versus the correct next character. We track training and test losses per shard and epoch.
@@ III.1 Traditional Reservoir Computing
Transformer Hidden Size $d_{h}=64$, Heads $h=4$, Layers $L=4$ $15067$
Hidden Size $d_{h}=72$, Heads $h=8$, Layers $L=8$ $30299$
Hidden Size $d_{h}=128$, Heads $h=8$, Layers $L=8$ $45083$
Hidden Size $d_{h}=356$, Heads $h=8$, Layers $L=8$ $105275$
Hidden Size $d_{h}=256$, Heads $h=16$, Layers $L=16$ $155803$
Reservoir Reservoir Size $N=250$ $14809$
Reservoir Size $N=500$ $29559$
Reservoir Size $N=750$ $44309$
Reservoir Size $N=1750$ $103309$
Reservoir Size $N=2600$ $153459$
Att Reservoir Reservoir Size $N=75$, Hidden Size $H=13$ $15464$
Reservoir Size $N=75$, Hidden Size $H=19$ $31124$
Reservoir Size $N=100$, Hidden Size $H=20$ $45259$
Reservoir Size $N=150$, Hidden Size $H=25$ $102809$
Reservoir Size $N=160$, Hidden Size $H=30$ $155459$
@@ IV Architectures
Input Text to be, or not clear born have the poor my turn with my soul. romeo what when queen let this shall let the sbatter did the death, well the part, that your grief is turned, we hard me, like on an i when a true, and her-promain the commone to with you this lodged prithee to be, or not as a find. hele to where i wheress thieve hot may come, and you the sease, thy agains for your gon. pring thy worney that light and he winder, eyes was from three like io be good reator, but have them beenes of be strengler. i have such a foglingers widdly, good time i, btund. to be, or not my could now richment. lodovico-ling, as yield make one in their if younger, farewell. duchall, then, my great, teater these is and that beauty brow man, for here is the lord of play birthing on the had your and, but i’ll gench i heart that this mans colors well TransformerReservoirAERCVector Embedding
@@ VI Methodology
We divide the Shakespeare corpus into $6$ shards. Five shards serve as training data, and one shard is used for testing. We sample sequences of length $32$, i.e. 32 characters as input, and train on a batch size of $1024$ with Adam at a learning rate $1\times 10^{-4}$. The traditional reservoir, the attention-enhanced reservoir, and the transformer models are trained entirely by backpropagation in a consistent framework. We measure cross-entropy on the test shard as our main performance metric. We also assess closed-loop generation by feeding the predicted character back as input, measuring distributional metrics and repetitive patterns. Our parameter sweep includes varying the reservoir size or the neural network size in the attention reservoir, as well as varying the hidden dimension $d$ and number of layers and heads $L$ in the transformer. To save compute power, we first compute all the reservoir responses in one shard and keep them in memory, after which a few epochs on this shard are run before the next shard is processed. This procedure is repeated multiple times for all shards.
@@ VIII Conclusion
We presented a unified framework for comparing classical reservoir computing (with a static linear readout), an attention-enhanced reservoir (with a dynamic readout), and transformer architectures on a character-level Shakespeare corpus. By systematically varying model sizes and measuring cross-entropy, N-gram Overlap, and time costs, we revealed how each architecture scales in performance and resource usage. Transformers excel overall, solidifying their state-of-the-art machine learning status, though they are more demanding in terms of computation time due to the quadratic complexity of the sequence length. Traditional reservoirs, while limited in their ability to capture complex dependencies, show surprising good results for their simple buildup. Attention-enhanced reservoirs strike a balance between efficiency and performance, outperforming traditional reservoirs and even getting close to the performance of Transformer models while having a negotiable complexity depending on the sequence length.

### v2

@@ I Introduction
[…]
@@ II Task Formulation and Dataset
We tackle a character-level next-character prediction task. Let $\mathbf{X}=\{x_{1},\dots,x_{T}\}$ represent a fixed-length input sequence of characters, and let $y_{T+1}$ be the next character to predict. We treat the problem as a multi-class classification over the vocabulary $\mathcal{V}=59$. The dataset is a 9-million-character corpus of Shakespeare text inspired by recent minimalist transformer implementations such as Karpathy’s NanoGPT [22], which we split into 666 shards: 555 shards for training and 111 shard for testing. Each shard is processed at the character level, forming sequences of length 323232 for the input, plus the next character as the target. Before training the text corpus is preprocessed by lowercasing all letters to achieve a smaller vocabulary $\mathcal{V}=59$. Training optimizes a standard cross-entropy loss over the predicted logits versus the correct next character. We track training and test losses per shard and epoch.
@@ III.1 Traditional Reservoir Computing
Transformer Hidden Size $d_{h}=64$, Heads $h=4$, Layers $L=4$ 150671506715067
Hidden Size $d_{h}=72$, Heads $h=8$, Layers $L=8$ 302993029930299
Hidden Size $d_{h}=128$, Heads $h=8$, Layers $L=8$ 450834508345083
Hidden Size $d_{h}=356$, Heads $h=8$, Layers $L=8$ 105275105275105275
Hidden Size $d_{h}=256$, Heads $h=16$, Layers $L=16$ 155803155803155803
Reservoir Reservoir Size $N=250$ 148091480914809
Reservoir Size $N=500$ 295592955929559
Reservoir Size $N=750$ 443094430944309
Reservoir Size $N=1750$ 103309103309103309
Reservoir Size $N=2600$ 153459153459153459
Att Reservoir Reservoir Size $N=75$, Hidden Size $H=13$ 154641546415464
Reservoir Size $N=75$, Hidden Size $H=19$ 311243112431124
Reservoir Size $N=100$, Hidden Size $H=20$ 452594525945259
Reservoir Size $N=150$, Hidden Size $H=25$ 102809102809102809
Reservoir Size $N=160$, Hidden Size $H=30$ 155459155459155459
@@ IV Architectures
to be, or not clear born have the poor my turn with my soul. romeo what when queen let this shall let the sbatter did the death, well the part, that your grief is turned, we hard me, like on an i when a true, and her-promain the commone to with you this lodged prithee to be, or not as a find. hele to where i wheress thieve hot may come, and you the sease, thy agains for your gon. pring thy worney that light and he winder, eyes was from three like io be good reator, but have them beenes of be strengler. i have such a foglingers widdly, good time i, btund. to be, or not my could now richment. lodovico-ling, as yield make one in their if younger, farewell. duchall, then, my great, teater these is and that beauty brow man, for here is the lord of play birthing on the had your and, but i’ll gench i heart that this mans colors well TransformerReservoirAERC
@@ VI Methodology
We divide the Shakespeare corpus into 666 shards. Five shards serve as training data, and one shard is used for testing. We sample sequences of length 323232, i.e. 32 characters as input, and train on a batch size of 102410241024 with Adam at a learning rate $1\times 10^{-4}$. The traditional reservoir, the attention-enhanced reservoir, and the transformer models are trained entirely by backpropagation in a consistent framework. We measure cross-entropy on the test shard as our main performance metric. We also assess closed-loop generation by feeding the predicted character back as input, measuring distributional metrics and repetitive patterns. Our parameter sweep includes varying the reservoir size or the neural network size in the attention reservoir, as well as varying the hidden dimension $d$ and number of layers and heads $L$ in the transformer. To save compute power, we first compute all the reservoir responses in one shard and keep them in memory, after which a few epochs on this shard are run before the next shard is processed. This procedure is repeated multiple times for all shards.
@@ VIII Conclusion
We presented a unified framework for comparing classical reservoir computing (with a static linear readout), an attention-enhanced reservoir (with a dynamic readout), and transformer architectures on a character-level Shakespeare corpus. By systematically varying model sizes and measuring cross-entropy, N-gram Overlap, and time costs, we revealed how each architecture scales in performance and resource usage. Transformers excel overall, solidifying their state-of-the-art machine learning status, though they are more demanding in terms of computation time due to the quadratic complexity of the sequence length. Traditional reservoirs, while limited in their ability to capture complex dependencies, show surprising good results for their simple buildup. Attention-enhanced reservoirs strike a balance between efficiency and performance, outperforming traditional reservoirs and even getting close to the performance of Transformer models while having a negotiable complexity depending on the sequence length. The code used for this study is available at https://github.com/fekoester/Shakespeare_Res.